OpenStudy (anonymous):

establish the identity. sin[x+ pi/2]=cos x

OpenStudy (anonymous):

use formula for sin(A + B) = sin A cos B + sin B cos A subsitute x for A; substitute pi/2 for B

OpenStudy (haseeb96):

pi\2 = 90 degree so sin(x+90)=cosx taking left.hand.side sin(a+b)=sinacosb+cosasinb sin(x+90)= sinx cos 90 + cos x sin 90 since , sin 90=1 and cos 90=0 sin(x+90)=sinx(0) + cos x (1) sin(x+90)= 0+cos x sin(x+90)= cos x hence proved

OpenStudy (anonymous):

got it out thanks you guys!